Reports


 

Laser system overview.pdf  , bailleux 2009

 

Lasers in used :

- Diode pump laser DP2, x2, 447nm (1mJ/pulse)

- Ti:S laser pumped by YLF laser, 800 nm (1.48mJ/pulse)

- Diode laser DTL-419QT, 527nm (200 uJ/pulse)

 

Main specifications of the system with the DP2-2:

Kejun_nss12_laser_2.pdf ; Calor_Liyuan DP brochure photonics

 

The 76,000 lead tungstate (PbWO4) crystals in the CMS electromagnetic calorimeter (ECAL) are radiation hard to high integrated dose, but suffer from dose rate dependent radiation damage. The light Higgs discovery potential is directly related to the energy resolution of the ECAL at LHC, and thus requires precision calibration. The ability to calibrate the crystal calorimeter precisely will be a major factor in determining its ultimate performance. The following 3 aspects are considered: precalibration in a test beam, in situ calibration using physics events and a precise and stable light monitoring system (laser).

 

2 lasers: one for the crystals ( 447nm) and one for electronic (800 nm). Laser should have enough power to generate  light equivalent to several hundred GeV in a few hundreds crystal at time. The light will be transported to the crystals via optical fibers and through 3 levels of distribution. The light from each pulse will be normalized using Si PN photodiodes at each of the  levels of distribution.

 

Laser for crystals: inject light pulses into each individual PbWO4 crystal to measure the optical

transmission  near the scintillation spectrum peak (~500nm), and occasionally, as a cross check (700nm). 

The red laser 800nm is used time to time to study the wavelength dependency of the transparency corrections.

 

3 levels of distribution: a light source and high-level distribution system which sends pulses to a

selected calorimeter element (1/2 barrel supermodule or endcap), and a 2-level distribution

system mounted on each calorimeter which delivers the pulses to the crystals.

 

The system is designed to continuously monitor the calorimeter in one of the 2 operational modes:

- Continuous in-fill mode (during 3 us gaps every 88,924 u sin the LHC beam),

- Stand-alone monitoring (outside LHC fills to follow the recovery of the PbWO4 crystals)

 

To meet the stringent requirements on the light monitoring precision and stability a new commercial diode-pumped blue laser ("DP2-447") has been commissioned and installed at CERN for the 2012 operation of the CMS ECAL (TiS laser before) The laser unit has a simple structure and is expected to be more reliable than the existing lamp-pumped lasers used by the monitoring system. The stability of critical quantities such as the intensity, width and timing, is better than that of the lamp-pumped lasers.

 

 

2018

integration forum, 15 Jan                                     laser_room_15Jan2018.pdf

 

2016

 

DPG Amina, end of year                                       161212_AZ_ECALDPGWorkshop.pdf

Status after Liyuan intervention 29March:              Laser status TC Avril 2016.pptx

ECAL DPG March:                                             MarcD_DPG_160329.pdf    intercalibration laser/VFE , temp effect from laser room..

                                    

 

2015

 

ECAL review 2-3Feb                                            Marc_ecalReviewFeb2015.pdf  ;  P_Gras_ecalReviewFeb2015.pdf

Calibration Workshop October                              calibration_Amina_oct2015.pdf

laser status Oct                                                       LasersOct2015.pptx

 

2014

 

 

OM 26 Nov                                                          pgras-20141124.pdf       ;     JLF_OM_26112014_R.pdf

MoCA 24 Nov                                                     MD_MoCa_141120.pdf

laser status cms week Oct                                      Laser status Oct 2014.pdf

OM 16april                                                           DP2_commision_GR.pdf       

MoCA August  (new monitorning box)                   MD_Moca_140828.pdf                                                      

WGM meeting 9 April (CMS wedn)                      CMS_wedn_meeting_WGM.pdf

OM meeting March, trigger effect on amplitude:     https://indico.cern.ch/event/306670/

Report before 2nd DP2 installation (DPG March)  DP2_12-658_summary.pdf

pcs installation:                                                     

OM meeting January:                                             https://indico.cern.ch/event/298179/

 

2013

 

laser_system_summary_3.pdf   Renyuan, December

ECAL_GRIN.pdf, Nov.

DP2_12-658_output_power_10-11-2013.pdf    2nd DP2 at Caltech after RH installation at Photonics.
Update on the DP2-447s_Oct-11-2013.pdf

Adi_ECALDAYS24Sept.pdf

report/reports2013/Update on the DP2-447s_Sept-23.pdf

Update on DP2-447_Caltech_July2013.pdf

Laser status at Caltech_2.pdf

laser_system_summary_1_April.pdf, Renyuan CMS week

diode_pump_laser, Kejun, ECAL conference committee , end 2012

Laser setup with the 2 Photonics .pdf

laser_room2013.pdf

power measurement 18Fev2013.pptx

2012

 

cms week Adi  Dec2012_Adi.pdf

OM Meeting 19Sept:  ZHU_OM_19Sept2012.pdf;     liyuan_OM_19Sept.pdf PowerDP2_OM_19Sept.pdf

power measurement 3Sept.pdf

power measurement 2AUGUST.pdf 

Tommaso_OM_1August.pdf

power measurement 11July.pdf

nss12_dp2_summary-1.pdf                Kejun, IEEE-NSS-MIC-RTDS2012 USA

OM_meeting_11July2012.pdf            OM meeting 11 July   power measurement 11July.pdf

Liyuan_CMS_week_June.pdf            https://indico.cern.ch/conferenceDisplay.py?confId=197823

OM_meeting_20June2012.pdf           OM meeting 20 June

power measurement 11 June.pdf         OM meeting 13 June

Calor_June_Liyuan                  agenda

PLOTS_LASERS_AprilMay.pdf

Incident history report.pptx      2012 list of incidents

Easter Incident Caltech group meeting.pptx     (wrong date, not easter but 1 May)     https://indico.cern.ch/conferenceDisplay.py?confId=168368

Power measurement ,                           OM meeting 2 May

Marc_12April.pdf

 

22 March monitoring meeting:

Kejun_DP2-447_Measurement_Caltech_March2012.pdf

Liyuan_DP2-447_Measurement_Caltech_March2012.pdf

Philippe_Control of new lasers.pdf

 

ZHU_CMS_Feb2012_without_Photonics_schematics.pdf

Marc_ECAL_FEb2012.pdf            New laser DP optical path, planning to do 2012.

Photonics_quote_440nm.pdf

Final layout UO 213.pdf

2011

 

D.Barney_newLaser_CMSweek.pdf

ZHU CMS week Nov.pdf

https://cms-docdb.cern.ch/cgi-bin/DocDB/ShowDocument?docid=5662

Summary laser choice.pptx            Dave B.

laser1_11Oct2011.pdf                    Laser1 problem 6-11 Oct (end of tuning 11 Oct)

wolfram_110929_Laser.pdf            Ecal days

ZHU_CaltehMeeting.pdf                Caltech meeting https://indico.cern.ch/conferenceDisplay.py?confId=119886

Lasers_27Sept2011.pdf                 Caltech and OM meeting. Matacq plots over 96hrs

Slow_Sept2011.pptx

lasers status August11.pdf

laser status 20July.pdf                     OM meeting

ecal_federico_matacq.pdf               CMS week

Laser_committee_cmsWeek.pdf

laser status 8JUNE.pdf                    OM meeting   

laser status 1JUNE.pdf                    OM meeting

LASER_UPGRADE-1_committee.pdf

Summary_Meeting_1_ECAL_Laser_TF.docx

optical_switch_cleanning.pdf

 

2010

 

ryz_101207_laser_replacement.pdf

CMS_week_march.pdf

Julie_M_29April2010.pdf

Federico_Ferri_Calor May 2010.pdf

Laser_EcalDays_Sep10.pdf

Julie_CMS_Nov2010.pdf

Laser_plots_2010.pptx

 

 

2009

 

LaserSystem review_15 Oct.pptx    Calibration and monitoring review

Dim_Kejun

Laser summary_3 August 09.docx

Laser Electronic Review 2009.pptx    Werner, electronics review part1 (doc)

Laser MO_March2009.pptx    US CMS meeting, CMS week

YLF laser upgrade.pdf    Diode PUMP

Laser tasks_CMS week March09.doc CMS week (for Werner)

Adi Milano Jan2009

B-field_equipment_1.xlsx

 

2008

 

B-field_onlaser_1.pdf                          * modify presentation to a report version

LASER under CMS field Nov.pdf          *Ecal pt5 commissioning meeting Nov

ECALDays_Laser_Monitoring_July2008_Adi.pdf

ECAL_Laser_cooling_v2[1].pdf

ryz_080422_award.pdf

Software feedback IEEE.pdf

08_tns_monitor.pdf

LASER_ecal_April_2008.pdf

P.Gras CMSweek.pdf

 

2007

 

Kejun_071029_N09-4_mon.pdf    Laser generale + feedback

Adi LaserPerformance2007.pdf

TB_2007-03-22_JV_PulseWidth_v2.pdf (J.Veverka)

Adi_March07 Rome.pdf

Kejun laser 7MArch2007

monitor2 Saclay, CMS internal note

Adi Data processing May2007.pdf

 

2006

 

nss06_laser_tns1.pdf                           (IEEE NSS at San Diego)

nss06_laser_1.pdf                                (Software feedback control)

Laser feedback Test Beam meeting 19July2006   

 

2005

 

cmspaper_2005.pdf

CMS_Conference_report 23Feb2005     (IEEE)

04_tns_monitor.pdf, IEEE Trans. Nucl. Sci.Vol. 52 No. 4 (2005) 1123--1130 (August 2005)

Adi_Nov2005 Pulsewidth.pdf

 

F http://agenda.cern.ch/fullAgenda.php?ida=a055740

LaserWorshop_David_16092005.pdf

LaserWorkshop_Adi_16092005[1].pdf

Adi_Test_beam_Sept05

Adi_VillaOlmo_Oct05.pdf

 

2004

 

CMS_week_Adi_June2004.pdf

David, Test beam _ 01 Sept 04.pdf, or web page

Adi_TB_Oct2004.pdf

CleanRoom_project.ppt, 28/Nov/04 (Alain Herve, Patrick Ingenito, Andrea Gaddi)

Ryz_041028_usc55.pdf,28/Nov/04 (Alain Herve, Patrick Ingenito, Andrea Gaddi)

Adi_TestBeam_2Nov2004

 

2003

 

ZHU_monitoring.pdf

Monitor_ins_2003.pdf                            (Installation Note, CMS IN 2003-045)

Test beam 15Nov2002.ppt

Test beam 25March2003.ppt

David_2003.ppt

Test beam meeting 090903.pdf

Adi_Milan2003.pdf